# An Axiomatic Risk-Reward Framework for Sustainable Investing

### v3

@@ 3 ESG-coherent risk measures
We begin with the axioms that define a coherent risk measure (Artzner et al.,, 1999). These axioms make it possible to identify measures with desirable properties, assisting both investors and regulators. Consider a convex set $\mathcal{X}\subseteq\mathcal{L}_{p}(\Omega,\mathcal{F},P)$ of real-valued random variables $r_{T}$ which are defined on a probability space $(\Omega,\mathcal{F},P)$, have finite $p$-moments ($p\geq 1$), and are indistinguishable up to a set of $P$-measure zero. We assume that the random variables represent the cumulative return over time T or the payoff at time T of an asset. The functional $\rho(r_{T}):\mathcal{X}\rightarrow\mathbb{R}\cup\{+\infty\}$ is a coherent risk measure if it satisfies the following properties:
@@
Examples of coherent risk measures are AVaR and expectile; but VaR and standard deviation are not coherent risk measures.101010 Alternative axiomizations of risk measures, such as convex risk measures (Föllmer and Schied,, 2002) and regular risk measures (Rockafellar and Uryasev,, 2013) have been proposed since the work of (Artzner et al.,, 1999).
The definition of a coherent risk measure is based upon a univariate random variable, and on the idea that the risk is a function only of the returns of the asset (consistent with the assumption that an investor is only interested in the monetary outcome of their position). In some contexts, the monetary outcome does not fully characterize the risk. Examples include: portfolios in two countries having floating exchange rates whose payoffs in the different currencies are not perfectly substitutable (the Siegel paradox, see Black,, 1989); the fact that various maturities for interest rate products are not perfect substitutes (failure of the pure expectation hypothesis); and cases in which it is difficult to attribute a monetary equivalent to various dimensions of risk, such as environmental or health risks. The same principle can be used for the ESG-oriented investor, whose risk depends on both the monetary return of a financial position and its sustainability, represented in our analysis by its ESG score. Such an approach allows an investor to deal with trade-offs between different goals that, although they are very different in nature, have to be taken into account in the investment process.
@@ 3.1 ESG-coherent risk measures — axiomatic definition
We introduce bivariate risk measures and apply them to the financial performance and sustainability of a position.111111 As discussed in Section 2, we use periodic returns for the monetary component and ESG scores (which represent the accumulated sustainability of the investment) for the sustainability component. The definition could be extended to alternative specifications that can be represented using two random variables with a joint distribution such that an investor has a preference for both higher $r$ and higher ESG{ESG} values (e.g., using the final net worth and the accumulated sustainability multiplied by the initial value of the position). Consider a convex set $\mathcal{X}_{2}$ of random vectors $X_{T}=[r_{T},\;ESG_{T}]^{\prime}$, defined on a probability space $(\Omega,\mathcal{F},P)$ with values in $\mathbb{R}^{2}$. We use the short-hand notation $\mathcal{X}_{2}=\mathcal{L}_{p}(\Omega,\mathcal{F},P;\mathbb{R}^{2})$ for the space of random vectors with two components and finite $p$-th moments, which are indistinguishable on sets of $P$-measure zero. Here $p\in[1,\infty]$. Since individual investors may have different attitude towards ESG, to highlights the trade-off between the monetary and sustainability risk components, we use the parameter $\lambda\in[0,1]$, a scalar which represents an investor preference for the relative weighing between the monetary and ESG components of risk.
@@
We then provide the following definition:
@@
We can readily show that an ESG-coherent risk measure has the following property TI-M, that generalizes the translation invariance properties to the bivariate case.
@@
  • The last axiom LM-H defines the value of the risk measure for a constant quantity and introduces the parameter $\lambda$ that accounts for the ESG preference of investors. It enables the characterization of the risk of an ESG safe asset (SA), i.e. a position having constant values for both monetary and ESG components. This specification ensures that $\boldsymbol{\rho}_{\lambda}(\boldsymbol{1})=-1,\quad\forall\lambda\in[0,1]$, where $\boldsymbol{1}$ is the vector $1,1]^{\prime}$.
@@
This set of axioms is related to the work of Rüschendorf, (2006), Wei and Hu, (2014), and Chen and Hu, (2020), which define convex risk for portfolio vectors using scalar-valued functions. A key difference from these papers, is the introduction of the parameter $\lambda$ in the axiom LM-H. Other authors have proposed set-valued risk measures (Jouini et al.,, 2004; Hamel,, 2009; Hamel et al.,, 2013; Feinstein and Rudloff,, 2013). Such proposals start from the idea that, “risk is the amount of cash that needs to be added to a position to make it acceptable” (Artzner et al.,, 1999). Extending this reasoning to a multivariate setting with multiple assets, it is possible to make a position acceptable by adding any of several “safe” portfolios with deterministic payoff, each characterized by a different combination of assets with deterministic payoff. In such a framework, the risk measure is given by a combination of all safe portfolios that make the position acceptable. In the context of ESG investing, where each asset is evaluated on the basis of two dimensions (returns and sustainability), we can imagine a market with multiple ESG safe assets with different return and ESG values. The advantage of this approach is to provide a more complete assessment of the risk in relation to the multiple drivers of risk, and has nicer mathematical properties, but at the cost of greater complexity and the inability to directly rank positions. Our approach, to compute a scalar risk based on a bivariate random variable, can be seen as a special case of the set-valued risk measures, in which we consider only one specific ESG safe asset.
@@
We note that the definition of an ESG-coherent risk measure remains agnostic concerning the measurement of either the financial performance or the ESG score; the former can be measured in terms of the final wealth, profit and loss, or periodic returns (Artzner et al.,, 1999), and the latter can be computed according to multiple methodologies and aggregated over time following several approaches. The only requirement is that the investor must have a preference for both higher financial gain and higher ESG scores (hence, the monetary part must be expressed in terms such that gains are positive).
In an analogous manner, we can define ESG-coherent reward measures that extend the work of Rachev et al., (2008) (see Appendix A).
@@ 3.1.1 Dual representation
It is well known that coherent risk measures have a dual representation - the supremum of a certain expected value over a risk envelope (Ruszczyński and Shapiro,, 2006; Ang et al.,, 2018; Dentcheva and Ruszczyński,, 2024). For ESG-coherent risk measures, the dual representation is introduced in Proposition 2.
@@ 3.2 Hedging risk by investing in ESG safe assets
To provide a more complete understanding of ESG-coherent risk measures, it is useful to study how ESG-oriented investors can hedge a risky position by investing in an ESG safe asset. In a traditional univariate framework, a safe asset by definition has a deterministic payoff;121212 Note that a position can have a risk equal to zero and not be a safe asset. Similarly, an asset with a deterministic payoff (i.e., a safe asset) can have a risk that is different from zero. We can understand this point better by considering a position with a return $r_{T}$ and risk $\rho(r_{T})=m$. If $\rho(\cdot)$ is a coherent risk measure, by axiom TI we have $\rho(r_{T}+m)=0$. That is, a position with a return $r_{T}^{\prime}=r_{T}+m$ has zero risk (but its returns are not necessarily constant). its return is a constant $\text{RF}_{T}^{r}\in\mathbb{R}$. If we define the risk on a univariate random variable that represents returns, axioms TI and PO state that the risk of a portfolio composed of the safe asset and a risky position $r_{T}$ is
@@
That is, the risky position can be hedged by constructing a portfolio that contains the safe asset having weight $w^{*}$.131313 If the risk measure were defined in terms of final net worth rather than returns, to hedge a risky position with risk $m$, it would be necessary to add a cash position. For a broader discussion of the interpretation of the axioms expressed in terms of returns rather than the final net worth, see Rachev et al., (2011, Chapter 6).
For ESG-oriented investing, an ESG safe asset is a position having constant values for both monetary and ESG components. We can postulate the existence of several types of such ESG safe assets, distinguished by different combinations of constant ESG and monetary return.
@@
    where $\text{RF}_{T}^{r}\geq 0$. An example of this is a risk-free, zero-coupon bond issued by a governmental institution not associated with a specific ESG profile.141414 Rating agencies are starting to compute ESG scores for countries as well, although the criteria are different from those used to calculate companies’ scores. The identification of pure monetary and pure ESG safe assets will be a significant challenge for practitioners and scholars. An ESG investor can hedge the risk of a position $X_{T}$ by constructing a portfolio composed of the pure monetary safe asset in proportion $w_{\lambda}^{*}$ and investing the rest in the risky asset, with
@@
    There exists an ESG safe asset having a monetary return of –100%. We consider the special case of an ESG safe asset described by
    $\text{SA}_{CHARITY,T}:=\begin{bmatrix}-1\\ \text{RF}_{T}^{{ESG}}\end{bmatrix}$.
    An example of this is a donation to a non-profit organization that has a positive and constant ESG score. Such an asset produces a monetary return of $100\%$, and clearly it is not a relevant investment opportunity for a risk-return investor with $\lambda=0$. [^15] Assuming the absence of tax benefits associated to the donation, or other forms of monetary gains. On the other hand, for an ESG-oriented investor with $\lambda>0$ it could be rational to invest in such asset. The measured risk of such an asset is $\boldsymbol{\rho}_{\lambda}(\text{SA}_{CHARITY,T})=(1-\lambda)-\lambda\text{RF}_{T}^{{ESG}}$; that is, for an investor with a given $\lambda$ the risk is negative if $\lambda>\tilde{\lambda}=1/(1+\text{RF}_{T}^{{ESG}})$. For a $\lambda<\tilde{\lambda}$” investor, such an asset provides no opportunity to hedge a risky position; however for a $\lambda>\tilde{\lambda}$” investor, such an ESG safe asset provides a meaningful hedging tool through the donation of a wealth fraction:
@@
The constraint set defines a convex feasible subset of $\mathbb{R}^{n}$ for the investment allocations, $w_{i}$. The solution is not a diversified portfolio; only one ESG safe asset is selected. This follows from the fact that, due to LH-M and TI-M, the risk of a sum of ESG safe assets is the sum of their risk: $\boldsymbol{\rho}_{\lambda}(\delta\text{SA}_{i,T}+(1-\delta)\text{SA}_{j,T})=\delta\boldsymbol{\rho}_{\lambda}(\text{SA}_{i,T})+(1-\delta)\boldsymbol{\rho}_{\lambda}(\text{SA}_{j,T}),\;\delta\in[0,1]$. For any pair of ESG safe assets $i$ and $j$, the risk of their convex combination is always greater or equal than $\min(\boldsymbol{\rho}_{\lambda}(\text{SA}_{i}),\boldsymbol{\rho}_{\lambda}(\text{SA}_{j})$. Hence to minimize the risk it is always convenient to choose the ESG safe asset with the smallest risk.161616 An exception is when two or more ESG safe assets with exactly the same risk are available. In such a case, they are indistinguishable to the investor in terms of risk, and either can be chosen. Once the ESG safe asset with the lowest risk is identified, the problem then is the same as (10) where only one ESG safe asset was available. Note however that the choice of ESG safe asset is different for each investor, as the optimization is dependent on the investor’s value for $\lambda$.
@@
In general, the choice of the ESG safe asset is not influenced by the characteristics of the risk of the position, it is only influenced by the availability and price of ESG safe assets and the $\lambda$ preference of the investor.
@@ 3.3 Examples of ESG-coherent risk measures
After discussing ESG-coherent risk measures in general, we present two approaches for extending univariate risk measures to a bivariate setting. In particular, starting from a univariate coherent risk measure $\rho(r_{T})$, we identify the ESG-coherent risk measures $\text{ESG-}\boldsymbol{\rho}_{\lambda}(X_{T})$, and $\text{ESG-}\boldsymbol{\rho}_{\lambda}^{l}(X_{T})$. The measures encompass $\lambda\in[0,1]$ as a parameter, thus they are families of bivariate risk measures suitable for investors having differing ESG “inclinations”. In Section 3.3.1 we apply these approaches to the well-known coherent risk measure, average value at risk (AVaR), resulting in two versions of coherent ESG-AVaR risk measures. In Section 3.3.2 we apply these approaches to two non-coherent risk measures, variance and volatility, producing ESG extensions that are not ESG-coherent.

### v4

@@ 3 ESG-coherent risk measures
We recall the axioms that define a coherent risk measure before introducing the bivariate framework (Artzner et al., 1999). These axioms make it possible to identify measures with desirable properties, assisting both investors and regulators. Consider a convex set $\mathcal{X}\subseteq\mathcal{L}_{p}(\Omega,\mathcal{F},P)$ of real-valued random variables $r_{T}$ which are defined on a probability space $(\Omega,\mathcal{F},P)$, have finite $p$-moments ($p\geq 1$), and are indistinguishable up to a set of $P$-measure zero. We assume that the random variables represent the cumulative return over time T or the payoff at time T of an asset. The functional $\rho(r_{T}):\mathcal{X}\rightarrow\mathbb{R}\cup\{+\infty\}$ is a coherent risk measure if it satisfies the following properties:
@@
Examples of coherent risk measures are AVaR and expectile; but VaR and standard deviation are not coherent risk measures. [^9] Alternative axiomizations of risk measures, such as convex risk measures (Föllmer and Schied, 2002) and regular risk measures (Rockafellar and Uryasev, 2013) have been proposed since the work of (Artzner et al., 1999).
The definition of a coherent risk measure is based upon a univariate random variable, and on the idea that the risk is a function only of the returns of the asset (consistent with the assumption that an investor is only interested in the monetary outcome of their position). In some contexts, the monetary outcome does not fully characterize the risk. Examples include: portfolios in two countries having floating exchange rates whose payoffs in the different currencies are not perfectly substitutable (the Siegel paradox, see Black, 1989); the fact that various maturities for interest rate products are not perfect substitutes (failure of the pure expectation hypothesis); and cases in which it is difficult to attribute a monetary equivalent to various dimensions of risk, such as environmental or health risks. The same principle can be used for the ESG-oriented investor, whose risk depends on both the monetary return of a financial position and its sustainability, represented in our analysis by its ESG score, that is not priced in monetary units. Such an approach allows an investor to deal with trade-offs between different goals that, although they are very different in nature, have to be taken into account in the investment process.
@@ 3.1 ESG-coherent risk measures — axiomatic definition
We now extend the classical axioms of coherent risk to a bivariate setting, where the two components represent monetary and sustainability dimensions of the same position. [^10] As discussed in Section 2, we use periodic returns for the monetary component and ESG scores (which represent the accumulated sustainability of the investment) for the sustainability component. The definition could be extended to alternative specifications that can be represented using two random variables with a joint distribution such that an investor has a preference for both higher $r$ and higher ESG{ESG} values (e.g., using the final net worth and the accumulated sustainability multiplied by the initial value of the position). Consider a convex set $\mathcal{X}_{2}$ of random vectors $X_{T}=[r_{T},\;ESG_{T}]^{\prime}$, defined on a probability space $(\Omega,\mathcal{F},P)$ with values in $\mathbb{R}^{2}$. We use the short-hand notation $\mathcal{X}_{2}=\mathcal{L}_{p}(\Omega,\mathcal{F},P;\mathbb{R}^{2})$ for the space of random vectors with two components and finite $p$-th moments, which are indistinguishable on sets of $P$-measure zero. Here $p\in[1,\infty]$. Since individual investors may have different attitude towards ESG, to highlight the trade-off between the monetary and sustainability risk components, we use the parameter $\lambda\in[0,1]$, a scalar which represents an investor preference for the relative weighing between the monetary and ESG components of risk. Such parameter is investor-specific but fixed at evaluation time.
@@
The first three axioms directly extend the univariate case, while LH-M axiom specifies the risk of deterministic ESG-monetary bundles and encodes investor preferences. We then provide the following definition:
@@
It follows from the axioms that any ESG-coherent risk measure satisfies the following property (TI-M), which generalizes translation invariance to the bivariate case.
@@
  • The last axiom LM-H defines the value of the risk measure for a constant quantity and introduces the parameter $\lambda$ that accounts for the subjective preference of each investor between financial and ESG dimensions.
  • The axiom LM-H enables the characterization of the risk of an ESG safe asset (SA), i.e. a position having constant values for both monetary and ESG components. This specification ensures that $\boldsymbol{\rho}_{\lambda}(\boldsymbol{1})=-1,\;\forall\lambda\in[0,1]$, where $\boldsymbol{1}$ is the vector $1,1]^{\prime}$.
@@
This set of axioms is related to the work of Rüschendorf (2006), Wei and Hu (2014), and Chen and Hu (2020), which define convex risk for portfolio vectors using scalar-valued functions. A key difference from these papers, is the introduction of the parameter $\lambda$ in the axiom LM-H.
@@
Further insights can be gathered by considering the acceptance set characterization of univariate risk measures: as discussed in Artzner et al. (1999), in the univariate case risk can be interpreted as the minimum amount of cash that has to be added to a position to make it acceptable (i.e. to have risk smaller or equal than zero). Extending to a multivariate setting, we no longer have a single safe asset (cash), but rather several instruments with deterministic payoffs. More specifically, in the context of ESG investing, one can postulate the presence of multiple such instruments, each characterized by a different combination of return and ESG attributes. In this case, it is possible to make a risky position acceptable by adding specific amounts of the available safe assets. Indeed, an alternative set-valued multivariate extension discussed in the literature defines risk as the set of all the deterministic capital positions that added to a risky position make it acceptable. The advantages of the set-valued approach is to provide a more complete assessment of risk in relation to the multiple drivers, and to provide a general mathematical formulation, but at the cost of greater complexity and the inability to directly rank positions. We limit our analysis to the scalar-valued bivariate risk measures illustrated above, and we refer to Jouini et al. (2004); Hamel (2009); Hamel et al. (2013); Feinstein and Rudloff (2013) for a broader discussion on set-valued multivariate risk. The practical implication for ESG-oriented investors of the presence of multiple safe assets in the market, are further discussed in Section 3.2.
We note that the definition of an ESG-coherent risk measure remains agnostic concerning the measurement of either the financial performance or the ESG score; the former can be measured in terms of the final wealth, profit and loss, or periodic returns (Artzner et al., 1999), and the latter can be computed according to multiple methodologies and aggregated over time following several approaches. The only requirement is that the investor must have a preference for both higher financial gain and higher ESG scores (hence, the monetary part must be expressed in terms such that gains are positive).
In an analogous manner, we can define ESG-coherent reward measures that extend the work of Rachev et al. (2008) (see Appendix A).
@@ 3.1.1 Dual representation
It is well known that coherent risk measures have a dual representation - the supremum of a certain expected value over a risk envelope (Ruszczyński and Shapiro, 2006; Ang et al., 2018; Dentcheva and Ruszczyński, 2024). For ESG-coherent risk measures, the dual representation is introduced in Proposition 2.
@@ 3.2 Hedging risk by investing in ESG safe assets
To clarify both the structure and the practical use of ESG-coherent risk measures, we analyze how ESG-oriented investors may hedge a risky position by allocating capital to an ESG safe asset, better clarifying the interpretation of the translation invariance property in the bivariate setting. We will consider the case of an investor or an asset manager that is subject to the budget constraint, and thus to invest in an ESG safe asset needs to disinvest an equal amount from the risky asset, following standard treatments in the risk-measure literature (see, e.g. Rachev et al., 2011, Chapter 6.4.4). In a traditional univariate framework, a safe asset by definition has a deterministic payoff; [^11] Note that a position can have a risk equal to zero and not be a safe asset. Similarly, an asset with a deterministic payoff (i.e., a safe asset) can have a risk that is different from zero. We can understand this point better by considering a position with a return $r_{T}$ and risk $\rho(r_{T})=m$. If $\rho(\cdot)$ is a coherent risk measure, by axiom TI we have $\rho(r_{T}+m)=0$. That is, a position with a return $r_{T}^{\prime}=r_{T}+m$ has zero risk (but its returns are not necessarily constant). its return is a constant $\text{RF}_{T}^{r}\in\mathbb{R}$. If we define the risk on a univariate random variable that represents returns, axioms TI and PO state that the risk of a portfolio composed of the safe asset and a risky position $r_{T}$ is
@@
That is, the risky position can be hedged by constructing a portfolio that contains the safe asset having weight $w^{*}$. [^12] If the risk measure were defined in terms of final net worth rather than returns, to hedge a risky position with risk $m$, it would be necessary to add a cash position. For a broader discussion of the interpretation of the axioms expressed in terms of returns rather than the final net worth, see Rachev et al. (2011, Chapter 6).
In the ESG-oriented investing framework, an ESG safe asset is a position having constant values for both monetary and ESG components. For illustrative purposes, we can postulate the existence of several types of such ESG safe assets, distinguished by different combinations of constant ESG and monetary return.
@@
    where $\text{RF}_{T}^{r}\geq 0$. An example of this is a risk-free, zero-coupon bond issued by a governmental institution not associated with a specific ESG profile. [^13] Rating agencies are starting to compute ESG scores for countries as well, although the criteria are different from those used to calculate companies’ scores. The identification of pure monetary and pure ESG safe assets will be a significant challenge for practitioners and scholars. An ESG investor can hedge the risk of a position $X_{T}$ by constructing a portfolio composed of the pure monetary safe asset in proportion $w_{\lambda}^{*}$ and investing the rest in the risky asset, with
@@
    There exists an ESG safe asset having a monetary return of –100%. We consider the special case of an ESG safe asset represented by
    $\text{SA}_{GRANT,T}:=\begin{bmatrix}-1\\ \text{RF}_{T}^{{ESG}}\end{bmatrix}$.
    The financial component corresponds to a deterministic loss of the entire investment, while the ESG component is positive and constant. As a conceptual benchmark, this position may be interpreted as a grant to a charitable organization generating a fixed ESG contribution. Such an asset is clearly not a relevant investment opportunity for a risk-return investor with $\lambda=0$. [^14] Assuming the absence of tax benefits associated to the donation, or other forms of monetary gains. On the other hand, for an ESG-oriented investor with $\lambda>0$ it could be rational to invest in such asset. The measured risk of such an asset is $\boldsymbol{\rho}_{\lambda}(\text{SA}_{GRANT,T})=(1-\lambda)-\lambda\text{RF}_{T}^{{ESG}}$; that is, for an investor with a given $\lambda$ the risk is negative if $\lambda>\tilde{\lambda}=1/(1+\text{RF}_{T}^{{ESG}})$. For a $\lambda<\tilde{\lambda}$” investor, such an asset provides no opportunity to hedge a risky position; however for a $\lambda>\tilde{\lambda}$” investor, such an ESG safe asset provides a meaningful hedging tool through the donation of a wealth fraction:
@@
The constraint set defines a convex feasible subset of $\mathbb{R}^{n}$ for the investment allocations, $w_{i}$. The solution is not a diversified portfolio; only one ESG safe asset is selected. This follows from the fact that, due to LH-M and TI-M, the risk of a sum of ESG safe assets is the sum of their risk: $\boldsymbol{\rho}_{\lambda}(\delta\text{SA}_{i,T}+(1-\delta)\text{SA}_{j,T})=\delta\boldsymbol{\rho}_{\lambda}(\text{SA}_{i,T})+(1-\delta)\boldsymbol{\rho}_{\lambda}(\text{SA}_{j,T}),\;\delta\in[0,1]$. For any pair of ESG safe assets $i$ and $j$, the risk of their convex combination is always greater or equal than $\min(\boldsymbol{\rho}_{\lambda}(\text{SA}_{i}),\boldsymbol{\rho}_{\lambda}(\text{SA}_{j})$. Hence, risk minimization is achieved by selecting the ESG safe asset with minimal risk, and not by forming portfolios of safe assets. This uniqueness property, that is a direct consequence of ESG-coherence of $\boldsymbol{\rho}_{\lambda}$, parallels the univariate setting, where, in the presence of multiple safe assets, the optimal choice is the cheapest one. [^15] An exception is when two or more ESG safe assets with exactly the same risk are available. In such a case, they are indistinguishable to the investor in terms of risk, and either can be chosen. Once the ESG safe asset with the lowest risk is identified, the problem then is the same as (10) where only one ESG safe asset was available. Note however that the choice of ESG safe asset is different for each investor, as the optimization is dependent on the investor’s value for $\lambda$.
@@
In general, for risk minimization under a budget constraint, the choice of the ESG safe asset is not influenced by the characteristics of the risk of the position, it is only influenced by the availability and price of ESG safe assets and the $\lambda$ preference of the investor.
@@ 3.3 Examples of ESG-coherent risk measures
After discussing ESG-coherent risk measures in general, we present two example approaches for extending univariate risk measures to a bivariate setting. In particular, starting from a univariate coherent risk measure $\rho(r_{T})$, we identify the ESG-coherent risk measures $\text{ESG-}\boldsymbol{\rho}_{\lambda}(X_{T})$, and $\text{ESG-}\boldsymbol{\rho}_{\lambda}^{l}(X_{T})$. The measures encompass $\lambda\in[0,1]$ as a parameter, thus they are families of bivariate risk measures suitable for investors having differing ESG “inclinations”. In Section 3.3.1 we apply these approaches to the well-known coherent risk measure, average value at risk (AVaR), resulting in two versions of coherent ESG-AVaR risk measures. These examples extend naturally to the ESG formulation without changing their convex-analytic structure. In Section 3.3.2 we apply these approaches to two non-coherent risk measures, variance and volatility, producing ESG extensions that are not ESG-coherent.